10x + 5y = 52.5
30x – 7y = 80.5

Answer by mananth(16946)

10 x + 5 y = 52.5 .............1
30 x -7 y = 80.5 .............2
Eliminate y
multiply (1)by 7
Multiply (2) by 5
70 x 35 y = 367.5
150 x -35 y = 402.5
Add the two equations
220 x = 770
/ 220
x = 3.50
plug value of x in (1)
10 x + 5 y = 52.5
35 + 5 y = 52.5
5 y = 52.5 -35
5 y = 17.5
y = 3.50